The temperature on the thermometer reads 45 straight F presuperscript straight o . What is this temperature in straight C presuperscript straight o ?
Answer #1

given

Tf = 45 F (degrees Faurenheit)

equivakent temperature, TC = ?

(Tc - 0)/(100-0) = (TF - 32)/(212-32)

Tc/100 = (TF - 32)/180

TC = (5/9)*(TF - 32)

= (5/9)*(45 - 32)

= 7.22 C (degrees celsius)